If demand function is given as the following: 

Qz = 230 -2.75 Pz + 0.5 I + 1.2 Pm + 0.6A 

Where Qz is quantity of Good z sold, Pz is price of Good z per unit, I is per capita income, Pm is price of competitor and A is the amount of advertising spent. 

Current values:  Pz= RM 55 I= RM 9000    Pm= RM 50     A =RM 12,000

Please Determine the total revenue maximizing price and quantity


1
Expert's answer
2021-04-30T10:47:34-0400

After substitution of current values we get Qz=11990-2.75Pz. Total revenue is PQ=Pz(11990-2.75Pz). In order to find revenue maximizing price we should find the first derivative of this function and to solve the equation TR'=0. So, P(11990-2.75P)'=11990-5.5P=0; P=2180; Q=5995